OpenStudy (anonymous):

For the polynomial, list each real zero and its multiplicity. Determine whether the graph crosses or touches the x-axis at each x-intercept. F(X)= 3(x-4)(x-1)^4

OpenStudy (freckles):

do you know how to find when the polynomial f is zero?

OpenStudy (freckles):

here is a hint if you have a*b=0 then a=0 or b=0 or both=0 so you solve x-4=0 and (x-1)^4=0 also you can determine the multiplicity of the zero by looking at the power

OpenStudy (anonymous):

So you just plug 0 in where x is?

OpenStudy (freckles):

no

OpenStudy (freckles):

you plug in 0 where y is or where f(x) is..

OpenStudy (anonymous):

But there is no Y in the equation?

OpenStudy (freckles):

y=f(x) this is why I said or where f(x) is

OpenStudy (freckles):

To find the x-intercepts you set y to 0 To find the y-intercept you set x to 0 --- f(x)=y

OpenStudy (freckles):

\[y=3(x-4)(x-1)^4 \\ 0=3(x-4)(x-1)^4\]

OpenStudy (anonymous):

so you then solve for x?

OpenStudy (freckles):

yes if you want to find the x-intercepts

OpenStudy (anonymous):

I'm still confused on how to fine the answer for what I want to find???

OpenStudy (freckles):

did you read what I first said.. "here is a hint if you have a*b=0 then a=0 or b=0 or both=0 so you solve x-4=0 and (x-1)^4=0 also you can determine the multiplicity of the zero by looking at the power "

OpenStudy (anonymous):

Yes, I read that...but (AGAIN) I am still confused...I don't understand what you mean??

OpenStudy (freckles):

so you are unable to solve x-4=0? and also unable to solve (x-1)^4=0?

OpenStudy (anonymous):

the first one would be x=4 I don't know how to do the second one...

OpenStudy (freckles):

do you know 0^4 is 0?

OpenStudy (freckles):

or you can think it like this (x-1)^4=0 is equivalent to (x-1)(x-1)(x-1)(x-1)=0

OpenStudy (freckles):

and again using that if you have a*b=0 then a=0 or b=0 or both=0

OpenStudy (freckles):

you have x-1=0 or x-1=0 or x-1=0 or x-1=0

OpenStudy (freckles):

can you solve x-1=0?

OpenStudy (anonymous):

x=1

OpenStudy (freckles):

right and you have that equation 4 times because of that power of 4 which tells you have x=1 multiplicity 4

OpenStudy (anonymous):

ok, so does it touch or cross the x-axis?

OpenStudy (freckles):

you can look at a graph but zeros associated with even multiplicities means you will have a touch and zeros associated with odd multiplicities means you will have a cross